OpenStudy (anonymous):

Let f(x) = x2 + 2x - 1 and g(x) = 2x - 4. Find 2f(x) - 3g(x) 2x^2-2x-14 I'm curious to know if I did this right?

OpenStudy (anonymous):

re look at your calculations again, you are close but just off by last term

OpenStudy (anonymous):

Like Juanita said, check the negative sign on your last term

OpenStudy (tkhunny):

If only you would have shown intermediate results. We could have told you where you did or didn't wander off. \(f(x) = x^{2} + 2x - 1\) \(g(x) = 2x - 4\) \(2f(x) = 2x^{2} + 4x - 2\) \(3g(x) = 6x - 12\) \(2f(x) - 3g(x) = (2x^{2} + 4x - 2) - (6x - 12) = 2x^{2} + 4x - 2 - 6x + 12\) \(2f(x) - 3g(x) = 2x^{2} + 4x - 6x - 2 + 12 = 2x^{2} - 2x + 10\) If I had to guess, and I do, since you didn't show your work, you did not properly utilize the Distributive Property right here \(-(6x - 12) = -6x + 12\).

OpenStudy (anonymous):

2x^2 + 4x -2 -6x + 12 2x^2 -2x -10 instead of -10 you add -2 and 12 as if they totaled 14

OpenStudy (anonymous):

I was told if f(x)-g(x) then the terms on g's side would become negative. I guess that is why I got 14

OpenStudy (anonymous):

g(x) = 2x - 4 3g(x) = 3(2x -4 ) = 6x -12 = -1( 6x - 12) = -6x + 12

OpenStudy (anonymous):

Yes each sign does change...because you are multiplying by a negative 1 (subtracting)

OpenStudy (anonymous):

When you have something like -g(x), you need to distribute the negative to the entire expression, example: g(x) = 2x - 4 -g(x) = -(2x - 4) but simplifies to: -g(x) = -2x + 4